OpenStudy (decarr432):

What is the range of the function f(x) = 3x − 12 for the domain {-2, 2} Please walk me through it medal and fan will be rewarded

OpenStudy (decarr432):

@zepdrix

OpenStudy (decarr432):

@Loser66 @mathmale

OpenStudy (loser66):

put x =-2 in f(x) , what do you get?

OpenStudy (decarr432):

I get -6x-12 right?

OpenStudy (decarr432):

f(x)=-6x-12

OpenStudy (loser66):

no more x

OpenStudy (loser66):

-6-12 =-18 , right?

OpenStudy (decarr432):

yeah okay

OpenStudy (decarr432):

so for second use 2?

OpenStudy (loser66):

how about replacing x =2?

OpenStudy (decarr432):

so f(x)=3(2)-12= 6-12= -6

OpenStudy (decarr432):

so answer is {-18,-6}?

OpenStudy (loser66):

yup

OpenStudy (loser66):

but not the curly bracket.

OpenStudy (decarr432):

That is so simply lol my teacher was trying to teach us this complicated crap

OpenStudy (loser66):

It should be [-18,-6]

OpenStudy (decarr432):

Oh for the answer choices it shows { not [

OpenStudy (loser66):

ok, so, just take it.

OpenStudy (decarr432):

Thanks mind helping me with a few more?

OpenStudy (loser66):

not sure, just post